Upon impact, bicycle helmets compress, thus lowering the potentially dangerous acceleration experienced by the head. A new kind of helmet uses an airbag that deploys from a pouch worn around the rider's neck. In tests, a headform wearing the inflated airbag is dropped onto a rigid platform; the speed just before impact is 6.0 m/sm/s. Upon impact, the bag compresses its full 12.0 cmcm thickness, slowing the headform to rest.

Part A

What is the acceleration, in gg's, experienced by the headform? (An acceleration greater than 60g60g is considered especially dangerous.)

Express your answer in units of the magnitude gg of the free-fall acceleration.
Answer #1

Using 3rd equation of motion,

v^2 = u^2 + 2as

0 = 6^2 + (2 x a x 0.12)

Acceleration, a = - 150 m/s^2

Acceleration, a = 150/9.8 = 15.3 g

Comment in case any doubt please rate my answer....